OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would multiply the following expression: ^12√x^5 times ^6√x (2 points) Part 2: Write the simplified answer in radical form. (2 points)

OpenStudy (anonymous):

\[\sqrt[12]{x^5}\times \sqrt[6]{}\]

jhonyy9 (jhonyy9):

the first term is equal with x on exponent 5/12 and the second is x on exponent 1/6 so when multiply two the same terms with exponents you need assuming the exponents so than will be x on exponent 5/12 + 1/6 = 7/12 so equal x on exponent 7/12